# DAViD: Modeling Dynamic Affordance of 3D Objects using Pre-trained Video Diffusion Models

### v1

@@ 4.2 Baselines and Metric
Following previous approaches on generating human motion with HOI, we evaluate the quality of our generated 4D HOI samples against FullBodyManipulation Dataset [49] in terms of (1) human motion, and (2) interaction quality. Specifically, we use the state-of-the-art method of 4D HOI sample generation, OMOMO [49] as a baseline of comparison. However, while the baseline generates 4D HOI samples with object motion guidance, allowing for comparison with ground truth, our approach uses only a 3D object as input, making it challenging to define the ground truth of our generated samples. To address this, we evaluate the interaction quality of the generated samples via contact patterns, which is widely used in the area of affordance learning.
Specifically, we aggregate the contacts shown in each frame of the generated 4D HOI sample to compute contact scores for human joints. The contact score for each point is computed as a possibility of contact during the HOI. We use mean average error (MAE) [101] and similarity score (SIM) [85] to compare contact scores against FullBodyManipulation Dataset [49]. For the human joints, as we use the output format of SMPL-X [69] (with the default hand pose) instead of the baseline’s SMPL [54], we only consider the contact possibility of wrist joints which have structurally the same position. To remove the bias of the left and right hand, we add the possibility of contact of both wrist joints to construct the final contact score. Additionally, to assess human motion quality, we measure Foot Sliding (FS) [30] and foot height $\text{H}_{\text{feet}}$ in centimeters (cm).
@@ 4.3 Qualitative Results
4D HOI Generation for Various Objects. Different from previous studies, we leverage the knowledge of a pre-trained video diffusion model, which allows to learn dynamic patterns of HOI for unbounded object categories objects. Fig. 4 visualizes the HOI motion for various objects generated by our DAViD. The frames are visualized in temporal order. The results show that our method effectively models the dynamic patterns naturally emerging in HOIs. In Fig. 4 (a) and Fig. 4 (b), we observe that the human pose and the relative position of the object are well modeled in relatively static HOI scenarios. In cases with more dynamic human motion, we find that various interaction patterns, including pulling--Fig. 4 (c), pushing--Fig. 4 (h), lifting--Fig. 4 (g), and holding--Fig. 4 (f) are well aligned with the typical use of objects of humans. Also, we can observe that our method can effectively capture complex human motions and their corresponding object motions, such as back and forth movement, as shown in Fig. 4 (e) and Fig. 4 (j). Importantly, unlike previous studies where motion translation typically resulted from walking, we observe that our model effectively captures cases of translation changes without foot movement, as shown in Fig. 4 (d) and Fig. 4 (h).
@@
Figure 5: Combining Concepts. Human motion concept learned by our LoRA can be combined with original knowledge of pre-trained MDM.
Figure 6: Ablation on SGS. SGS finds a smooth path that connects framewise plausible object poses at inference, increasing the plausibility of object motion.
Combining Multiple Concepts. As we model dynamic patterns of human motion using LoRA [33] on a pre-trained MDM [88], the existing knowledge of MDM [88] can be combined with the newly learned concepts of human motion. Fig. 5 shows the result of combining the directional concept from the pre-trained MDM [88] with the motion of using a motorcycle, which we learned through LoRA [33]. Specifically, we use text prompts such as “A person turns left riding a motorcycle”. Even without annotating direction knowledge during LoRA [33] tuning, we can combine our human motion concepts with the directional knowledge from the original MDM [88] to generate plausible HOI motion involving a motorcycle with change of direction.
@@
Ablation on Smoothness Guidance Sampling. To verify the efficacy of our SGS, we conduct an ablation study. Fig. 6 shows six continuous frames of HOI motion with a shopping cart, along with an overlay of consequent two frames. Without SGS, object poses are independently computed for the human pose, causing motion jitters between adjacent frames and reducing the plausibility of the motion, as shown in Fig. 6 (a). In contrast, with our SGS, the generated HOI motion shows relatively smooth as shown in Fig. 6 (b). This shows the importance of SGS as it finds a path that continuously connects plausible object poses for each frame, creating a motion that is both frame-wise plausible and temporally smooth.
@@
Ours 0.809 0.64 0.386 0.0654
OMOMO [49] 1.52 2.68 0.273 0.113
@@ 4.4 Quantitative Results
To verify the efficacy of our trained DAViD, we compare OMOMO and ours on the FullBodyManipulation Dataset [49]. We compare the quality of HOI motion generated by OMOMO, trained on the FullBodyManipulation Dataset with ours generated from DAViD, which is trained using synthetic 4D HOI samples. Specifically, we generate 50 HOI motions per category using DAViD, while OMOMO [49] generates HOI motions in the same quantity as the test dataset in the FullBodyManipulation Dataset [49]. For evaluating human quality, we compute the FS and Foot Height for each sample and then average them. As shown in Tab. 1, the quality of our human motion outperforms the baseline, indicating that our synthetic data is reliable enough to be comparable with real-world capture data. We demonstrate the effectiveness of our approach to learning from the synthetic data as it is easy to scale for various object categories. For evaluating interaction quality, we follow OMOMO [49] which focuses on comparing hand contact. We compute the hand contact score for both ours and baseline, and evaluate how similar these scores are to the hand contact score in the test set of FullBodyManipulation Dataset [49]. As shown in Tab. 1, the hand contact in our HOI motion accurately models the contact shown in the dataset than the baseline. According to the result, we demonstrate that our uplifting pipeline effectively decides spatial relations between humans and objects considering human-object contact, and consequently, DAViD well models the contact patterns.
Figure 7: Generating Multi-Object Interaction. Two human motion concepts learned by our LoRA is combined to generate multi-object interaction.
4.5 Application
As we learn human motion concepts using LoRA of the pre-trained MDM, it is possible to generate HOI motions with multi-object interactions by combining two different concepts. Specifically, we merge two separately trained LoRA parameters into the original network by a weighted summation, combining the two concepts into a single network. In practice, we assign the same weight value of 0.65 to each network parameter for summation. After merging the two LoRA models, we first generate human motion using the combined network, and then use this as a condition for the object pose diffusion model, obtaining each object’s pose sequence through SGS. Fig. 7 shows the result of combining two separately learned human motion concepts: one using a scooter and the other using an umbrella. We observe that the generated result shows plausible motion while satisfying the dynamic patterns of each object’s usage (e.g., holding for the umbrella, riding for the scooter). By generating multi-object interactions through a simple plug-and-play approach with LoRA layers, we demonstrate the advantages of using LoRA to learn human motion concepts in HOI.
@@ 5 Discussion
In this paper, we present a method to learn an affordance called Dynamic Affordance and our framework DAViD which models it, addressing that there exists not only static patterns (e.g., contact, spatial relation) but also dynamic patterns in HOIs. While obtaining a 4D HOI dataset including diverse objects is challenging, we observe that the video diffusion model has prior knowledge of the dynamic patterns for general objects. Compared to learning affordance from videos generated from text-to-video models, our pipeline, which first renders the 3D object to generate HOI images and then uses an image-to-video model, has the advantage of lifting affordances in video into 3D space. Using the 4D HOI samples generated through our lifting pipeline, we train DAViD, which consists of (1) the LoRA module of MDM for learning human motion patterns and (2) an object pose diffusion model with human pose guidance. We qualitatively show that the 4D HOI samples generated by DAViD effectively model the dynamic pattern of human pose and corresponding object pose during HOIs. Quantitative results show that our DAViD models human motion and interactions that are more similar to real-world HOI than the baseline. In addition, we demonstrate the advantage of using LoRA to model the dynamic patterns of using specific objects by generating HOI motion that include multi-object interactions.
@@ Acknowledgements
This work was supported by NAVER Webtoon, NRF grant funded by the Korean government (MSIT) (No. 2022R1A2C2092724, No. RS-2023-00218601), and IITP grant funded by the Korea government (MSIT) [No. RS-2024-00439854 , RS-2021-II211343]. H. Joo is the corresponding author.
@@ A.5 Network Architecture
Figure S.3: Additional Qualitative Results. We showcase additional results of our method. As shown in (a), our method effectively models the patterns during HOI. Furthermore, in (b), we present diverse samples generated from our DAViD, with each frame visualized in temporal order.
@@ A.6 Training Details
LoRA for MDM. For training the LoRA [33] layer in the pre-trained MDM [88], we create a dataset by extracting only the human motion from previously generated 4D HOI samples and processing it following HumanML3D [25]. The number of training samples varies by object category, ranging from $5$ to $15$, and we figure out that this amount is sufficient for learning the concept of human motion through LoRA [33]. During training, we freeze all other weights and train only the weights of the LoRA [33] layer. As our concepts are represented in the form of text, we prepare training text prompt of (1) simplified text and (2) tag. The simplified text is a modified version of the text prompt used earlier for 2D HOI image generation (Sec. A.2), formatted as “A person {action} {category},” while the tag represents the category of an object. For each training sample, one of the two text prompts is assigned for training. We train a total of 500 to 3000 steps (depending on categories) using the Adam [41] optimizer with a learning rate of $1\times 10^{-4}$ without decay.
Human Conditioned Object Pose Diffusion Model. For training human conditioned object pose diffusion model, we extract pairwise human pose and object pose from each frame of the 4D HOI Sample and use them as training data. The number of data samples used varies by object category, ranging from $765$ to $2,295$. For human poses, we find that using the SMPL-X [69] body pose parameters $\bm{\theta}\in\mathbb{R}^{21\times 3}$ produces qualitatively better results compared to using canonicalized point clouds used in GenPose [113]. To achieve this, we replaced the PointNet++ [73] encoder with 3-layer MLP. We train total of 1000 to 5000 steps (depending on categories) using the Adam [41] optimizer with a learning rate of $5\times 10^{-3}$ and a weight decay of 0.99.
@@ B.1 Additional Qualitative Results
We showcase additional qualitative results in Fig. S.3. In Fig. S.3(a), we show the results of training DAViD on additional object categories. As shown in the examples of the trumpet and cello, our method effectively models the relative position and orientation between human and the object, as well as the dynamic patterns, when playing instruments. Specifically, when using a trumpet, we observe a dynamic pattern where the hands hold the trumpet and keep it near the mouth while changing the head orientation. Similarly, when using a cello, we observe a dynamic pattern where the cello is positioned between the legs, and the right hand moves to play it. Fig. S.3(b) shows the results of generating various HOI motions using our trained LoRA [33]. Through the results of generating various HOI motions, we demonstrate that our LoRA [33] faithfully learns the dynamic patterns during HOI.
B.2 Scale of the Object

### v2

@@ 4.2 Baselines and Metric
Even our goal is to learn dynamic HOI patterns for various objects, which is slightly different from typical text-driven HOI motion synthesis [14, 70, 108, 112, 107, 103], we conduct quantitative comparisons on existing datasets for evaluation. Specifically, we compare DAViD against OMOMO [50] and CHOIS [51] on FullBodyManipulation Dataset [50], which provide available code and evaluate the quality of HOI, unlike other stuides [14, 70, 103] which focus on evaluating human quality solely. However, OMOMO [50] and CHOIS [51] generate human motion conditioned on object motion (or waypoints), making a direct comparison with ours challenging. To address this, we compare the HOI motion generated by baselines using object motion (or waypoints) condition with the HOI motion generated by DAViD using human motion condition. Each condition is extracted from the test dataset of FullBodyManipulation [50], enabling a comparison with the ground truth. As using different condition, we compare only the quality of HOI for the fairness, specifically the precision, recall, and F1 score of the contact, following baselines [50, 51]. As DAViD performs modeling of object poses during interaction, we evaluate by cropping each test dataset sequence from the start of the interaction (contact start) to the end (contact end). Additionally, to assess the human motion quality, we report the Foot Sliding (FS) and foot height (in centimeters) of the human motion generated from DAViD and the baselines, both trained on FullBodyManipulation Dataset [50].
@@ 4.3 Qualitative Results
4D HOI Generation for Various Objects. Different from previous studies, we leverage the pre-trained video diffusion model, which allows to learn dynamic patterns of HOI for unbounded object categories objects. Fig. 4 visualizes the HOI motion for various objects generated by DAViD. The frames are visualized in temporal order. The results show that our method effectively models the dynamic patterns in HOIs. In Fig. 4 (a), Fig. 4 (b), Fig. 4 (c) we observe that the human motion and the relative position of the object are well modeled in relatively static scenarios. In cases with more dynamic human motion, we find that various interaction patterns, including pulling--Fig. 4 (d), pushing--Fig. 4 (h), lifting--Fig. 4 (g), and holding--Fig. 4 (f) are well aligned with the typical use of objects of humans. Also, we can observe that our method can effectively capture complex human motions and their corresponding object motions, such as back and forth movement, as shown in Fig. 4 (e) and Fig. 4 (j). Importantly, unlike previous studies where motion translation typically resulted from walking, we observe that our model effectively captures cases of translation changes without foot movement, as shown in Fig. 4 (i).
@@
Figure 5: Combining Concepts. Our HOI motion concepts can be combined with original knowledge of pre-trained MDM, even for the multiple HOI motion concepts.
Figure 6: Ablation on Temporal Guidance Sampling. Our guidance during the inference stage reduces object jitter and non-contact interactions, enabling the generation of plausible HOI motion.
Combining Multiple HOI Concepts. As we model dynamic patterns of human motion using LoRA [31] on a pre-trained MDM [88], the existing knowledge of MDM [88] can be combined with the newly learned concepts of HOI. Fig. 5 (a) and Fig. 5 (b) show the novel HOI motion generated by combining the pre-trained MDM’s previously known jumping and walking concepts with the newly learned HOI concept; umbrella, and coffee. By simply adding LoRA tags along with the text prompt, we can combine the newly learned HOI concept with the existing motions known by pre-trained MDM to create a novel human motion. The corresponding object motion is then generated based on the human motion, forming the novel HOI motion.
This approach goes beyond combining a single HOI motion concept as shown in Fig. 5 (c), and allows to combine multiple HOI motion concepts which share a similar body structure. We merge two separately trained LoRA (suitcase, umbrella) parameters into the original network by a weighted summation, combining the two concepts into a single network. The object motions are then generated similarly based on the human motion.
@@
Ablation on Temporal Guidance Sampling. To verify the efficacy of our TGS, we conduct an ablation study. Fig. 6 shows six continuous frames of HOI motion with a suitcase, along with an overlay of objects in the subsequent two frames and a magnified view of the hand-object interaction. Without smoothness guidance, object poses are independently sampled for the human pose, causing motion jitters between adjacent frames and reducing the plausibility of the motion. Additionally, without contact guidance, fine details such as hand contact are lacking, resulting in floating objects. This shows the importance of TGS as it finds a path that continuously connects plausible object poses for each frame, creating a motion that is frame-wise plausible, temporally smooth, and contacted during interaction.
@@
CHOIS [51] 0.761 0.449 0.565 0.472 3.47
$\text{DAViD}_{\text{w/o contact guidance}}$ 0.923 0.154 0.264 0.234 2.30
$\text{DAViD}_{\text{Full}}$ 0.882 0.577 0.698
OMOMO [50] 0.783 0.547 0.644 0.255 2.58
$\text{DAViD}_{\text{w/o contact guidance}}$ 0.800 0.186 0.302 0.247 2.35
$\text{DAViD}_{\text{Full}}$ 0.828 0.558 0.667
MDM [88] N/A N/A N/A 0.260 2.48
@@
[…]
DAViD 30.9
$\text{DAViD}_{\text{w/o smoothness guidance}}$ $2.95\times 10^{3}$
FullBodyManip DB [50] 10.7
[…]
DAViD 1.30 8.90
$\text{DAViD}_{\text{w/o LoRA}}$ 2.93 8.59
Real  [24] 0.00160 9.48
Table 2: Quantitative Ablation. Our smoothness guidance reduces the jitter level of the object motion (left), and our LoRA module better preserves the pre-trained MDM knowledge (right).
@@ 4.4 Quantitative Results
Comparison with Baselines. To verify the efficacy of our model, we compare DAViD against OMOMO [50] and CHOIS [51] on the FullBodyManipulation Dataset [50]. Specifically, we measure the performance of DAViD on each of the test datasets of the FullBodyManipulation [50], selected from baselines. For the metrics about HOI quality, hand contact is considered when the average distance between the 25 hand joints of SMPL-X [69] and the object vertices is smaller than a threshold, 0.06. As shown in Tab. 1, DAViD outperforms the baselines for all contact metrics. This shows that DAViD generates plausible HOI motion, demonstrating the efficacy of our approach in modeling the object pose distribution from the human pose, and sampling object motion using TGS. Additionally, we report FS and foot height of the human motions generated by DAViD, where we generate the same number of samples as in the test dataset. We demonstrate that DAViD generates more natural human motions compared to the baselines, with lower FS and foot height, as shown in Tab. 1.
Ablation Study. We conduct a quantitative ablation study on our key design choices, TGS and LoRA [31]. As shown in Tab. 1, we find that our contact guidance in TGS is crucial for expressing detailed hand contact in HOI motions. Note that this is not obvious, as contact guidance automatically detects potential contact region and guides them to get closer, and not provide information on where the contact should occur. The result shows that the contact guidance in TGS performs fine refinement of object motion within a learned distribution. We also perform an ablation study on our smoothness guidance in TGS by generating object motion using human motion from FullBodyManipulation [50] and measure the jitter. We define jitter following MocapEvery [46] and compute it for object translation. As shown in Tab. 2 (left), our smoothness guidance in TGS significantly reduces the object jitter that occurs when sampling object poses framwise, bringing it to a level similar to the jitter found in real-world capture data.
We conduct an ablation study on LoRA to verify its ability to maintain the pre-trained model’s knowledge while extending the model’s space. We train both models for 1000 epochs on the FullBodyManipulation dataset [50] and compare the FID and diversity on HumanML3D dataset [24] (which MDM is trained on) for text-conditioned motion generation. As shown in Tab. 2 (right), LoRA enables the model to learn new concepts without significantly altering the motion distribution of the pre-trained model. We demonstrate that our design, which adds LoRA to MDM is important to preserve the knowledge of pre-trained model, allowing to combine various HOI concepts with pre-trained human motion.
@@ 5 Discussion
In this paper, we present DAViD, a method to learn Dynamic Affordance, addressing that there exists not only static patterns (e.g., contact, spatial relation) but also dynamic patterns in HOIs. While obtaining a 4D HOI dataset including diverse objects is challenging, we observe that the video diffusion model has prior knowledge of the dynamic patterns for general objects. Compared to learning affordance from videos generated from text-to-video models, our pipeline, which first renders the 3D object to generate HOI images and then uses an image-to-video model, has the advantage of lifting affordances in video into 3D space. Using the 4D HOI samples generated through our lifting pipeline, we train DAViD, which consists of (1) the LoRA module of MDM for learning human motion patterns and (2) an object pose diffusion model with human pose condition. We qualitatively show that the 4D HOI samples generated by DAViD effectively model the dynamic pattern of human pose and corresponding object pose during HOIs. Quantitative results show that our DAViD models human motion and interactions that are more similar to real-world HOI than the baseline. In addition, we demonstrate the advantage of using LoRA to model the dynamic patterns of using specific objects by generating HOI motion that include multi-object interactions.
@@ Acknowledgements
This work was supported by NAVER Webtoon, NRF grant funded by the Korean government (MSIT) (No. RS-2022-NR070498, No. RS-2023-00218601), and IITP grant funded by the Korea government (MSIT) [No. RS-2024-00439854, RS-2021-II211343]. H. Joo is the corresponding author.
@@ A.5 Network Architecture
Figure S.3: Additional Qualitative Results. We showcase additional results of our method. We present diverse samples generated from our DAViD, with each frame visualized in temporal order.
@@ A.6 Training Details
LoRA for MDM. For training the LoRA [31] layer in the pre-trained MDM [88], we create a dataset by extracting only the human motion from previously generated 4D HOI samples and processing it following HumanML3D [24]. The number of training samples varies by object category, ranging from $5$ to $50$, and we figure out that this amount is sufficient for learning the concept of human motion through LoRA [31]. During training, we freeze all other weights and train only the weights of the LoRA [31] layer. As our concepts are represented in the form of text, we use object category as a text prompt for training our LoRA. For motions with multiple modes (e.g., left and right hand-object interactions), the text prompt is modified by adding tags such as “left_” or “right_” before the main tag. We found that simply adding these additional tags gives controllability to model. We train a total of 500 to 3000 steps (depending on categories) using the Adam [39] optimizer with a learning rate of $1\times 10^{-4}$ without decay.
Human Conditioned Object Pose Diffusion Model. For training human conditioned object pose diffusion model, we extract pairwise human pose and object pose from each frame of the 4D HOI Sample and use them as training data. The number of data samples used varies by object category, ranging from $765$ to $7,650$. We train total of 1000 to 5000 steps (depending on categories) using the Adam [39] optimizer with a learning rate of $5\times 10^{-3}$ and a weight decay of 0.99.
@@ B.1 Additional Qualitative Results
We showcase additional qualitative results in Fig. S.3. In Fig. S.3(a), we show the results of training DAViD on additional object categories. Fig. S.3 shows the results of generating various HOI motions using our trained LoRA [31]. Through the results of generating various HOI motions, we demonstrate that our LoRA [31] faithfully learns the dynamic patterns during HOI.
B.2 Additional Quantitative Results
We conduct an additional quantitative evaluation on our method estimating object motion from the given video using PnP [19, 48] and 2D tracking [35, 34]. As shown in Tab. 3, our method outperforms the existing 6D pose estimation method, FoundationPose [101] in predicting 6D object poses in a single view at BEHAVE sequences via the commonly used ADD metric. As the FoundationPose [101] leverages depthmap from RGBD camera for 6D pose estimation, and our method only need RGB video, we use esimated depth for FoundationPose, which is from the metric depth model [4]. We also report the score of the best view for better assessment, as BEHAVE sequences contain object occlusion in a single view, which hinders estimating 6D object pose.
B.3 Scale of the Object
@@
[…]
Ours 43.40 64.96 85.52
$\text{FoundationPose}_{\text{w/ depth map estimation}}$ 24.67 29.91 40.89
Table 3: Object Pose Prediction against BEHAVE sequences.